OpenStudy (anonymous):

16 3/4 -3 7/8 can someone show me the way to figure this out by pencil

OpenStudy (anonymous):

start by combining the whole numbers... 16-3=13

OpenStudy (anonymous):

you are left with 3/4 - 7-8=? find the lowest common denominator (LCD)

OpenStudy (anonymous):

the LCD is 8 so you multiply the (3/4) by 2 and the withh give you 6/8

OpenStudy (anonymous):

now solve 6/8 - 7/8= -1/8

OpenStudy (anonymous):

13-1/8=12 7/8